OpenStudy (vera_ewing):

Find the zero of the function: f(x) = 5x - 3

OpenStudy (vera_ewing):

3/5 right? @Hero

hero (hero):

f(3/5)= 5( 3/5) -3= 0 correct

OpenStudy (vera_ewing):

Thanks.

hero (hero):

just showing you how to check your own answers.